# Regularity of semi-valuation rings and homotopy invariance of algebraic K-theory

### v2

@@ 1. Introduction
Semi-valuation rings have been introduced by Temkin [Tem11] and they occur as the stalks of his relative Riemann-Zariski spaces $\operatorname{RZ}_{Y}(X)$ which are locally ringed spaces associated with any separated morphism $f\colon Y\overset{}{\rightarrow}X$ between quasi-compact and quasi-separated (qcqs) schemes. Furthermore, semi-valuation rings are important since they are the stalks of discretely ringed adic spaces, see Remark 2.5. Exploiting that algebraic K-theory commutes with the formation of stalks, we deduce the following statement for the K-theory sheaves on this space (Definition 5.6).
@@
Let $f\colon Y\overset{}{\rightarrow}X$ be a separated morphism of finite type between qcqs schemes. Assume that $Y$ is of finite dimension, that all its stalks are stably coherent regular rings, and that the morphism $f\colon Y\overset{}{\rightarrow}X$ admits a compactification (e.g. $f$ is separated of finite type). Then the canonical maps
@@
Notation. Discrete categories are denoted by $\mathrm{upright}$ $\mathrm{letters}$ whereas genuine $\infty$-categories are denoted by $\mathbf{bold}$ $\mathbf{letters}$. For a ring $R$ we denote by $\operatorname{K}(R)$ its non-connective algebraic K-theory spectrum à la Blumberg-Gepner-Tabuada [BGT13, §9.1] which is an object of the $\infty$-category $\mathbf{Sp}$ of spectra [Lur17, 1.4.3.1]; its associated object in the homotopy category of spectra is equivalent to the K-theory spectrum of Thomason-Trobaugh [TT90, Def. 3.1]. In particular, the connective part $\operatorname{K}_{\geq 0}(R)$ is equivalent to Quillen’s K-theory [Qui73], see [TT90, Prop. 3.10].
@@ 2. Semi-valuation rings
For the convenience of the reader we recollect the definition and some basic facts about semi-valuation rings. All this is due to Temkin [Tem11].
@@
A semi-valuation ring is a ring $A^{+}$ with a valuation $|-|\colon A^{+}\overset{}{\rightarrow}\Gamma\cup\{0\}$ such that
@@
    every zero divisor of $A^{+}$ lies in the kernel of |||-|| - | and
@@
If $A^{+}$ is a semi-valuation ring and $\mathfrak{p}:=\ker(|-|)$, then the local ring $A^{+}_{\mathfrak{p}}$ is called its semi-fraction ring and the ring $A^{+}/\mathfrak{p}$ is a valuation ring. We call $\mathfrak{p}$ the valuative ideal of $A^{+}$.
@@
    A semi-valuation ring $A^{+}$ with an equivalence class of a valuation |||-|| - |.
@@
[…]
@@
In the sequel, we refer to a semi-valuation ring as a pair $(A^{+},\mathfrak{p})$ as in Lemma 2.2 and always assuming the square ()($\square$)( □ ) implicitly defined.
@@ 3. Coherence and regularity of semi-valuation rings
    $n$-regular if every finitely $n$-presented $A$-module has finite projective dimension.
@@
The implications of (i) and (ii) hold by design. The implication in (iii) follows since projective modules over local rings are free. The implication (vi)$\Rightarrow$(v) is trivial. If $A$ is coherent, every finitely presented $A$-module is pseudo-coherent and vice versa, hence (iv)$\Leftrightarrow$(v). The equivalence with (vii) goes by induction on the number of generators [Gla89, Thm. 6.2.1]. ∎
@@
The notion “1-regular” is called “regular” by Gersten [Ger74, Def. 1.3]. The notion “Glaz-regular” is called “regular” by Glaz [Gla89, ch. 6, §2]. Antieau-Mathew-Morrow [AMM22, §2] call a ring “weakly regular” if it has finite flat dimension (for them a regular ring is noetherian). For coherent rings, the notion “weakly regular” is equivalent to the notion “uniformly 1-regular” and it is stronger than the notions used by Gersten and Glaz. For arbitrary rings, the notion “regular” from Definition 3.2 seems to be most meaningful (at least to the author).
@@
[…]
$00=\operatorname{Tor}_{1}^{R}(M,R/\mathfrak{m})\overset{}{\longrightarrow}N/\mathfrak{m}N\overset{}{\longrightarrow}(R/\mathfrak{m})^{n}\overset{\cong}{\longrightarrow}M/\mathfrak{m}M\overset{}{\longrightarrow}$ 0
@@
be an exact sequence where $F_{n+1},\ldots,F_{0}$ are free. Setting $K_{i}:=\ker(f_{i})$ we get that
@@
${A^{+}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$${A\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{ev}_{0,0}}$${\mathbf{Z}_{p}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbf{Q}_{p}.}$
@@
[…]
@@
[…]
@@
where $\mathbf{RMod}(-)$ denotes the presentable, stable $\infty$-category of (derived) right modules, $\mathbf{Perf}(-)$ its full subcategory of perfect modules (which are precisely the compact objects), and “$\overset{\rightarrow}{\times}$” the lax pullback [Tam18, Def. 5]. The horizontal functors (which are induced by the base change functors) are fully faithful [LT19, 1.7] and the left vertical functor is the inclusion of compact objects by definition. The right horizontal functor identifies also with the inclusion of compact objects [Tam18, Prop. 13]. Now the claim follows since base change, and hence the horizontal functors, preserve filtered colimits [Tam18, Lem. 8 (iii)]. ∎
@@
We know that the ring $A^{+}$ is 2-regular by Lemma 3.7. If it was also coherent, then it would be Glaz-regular by Proposition 3.3, hence $A$ would be Glaz-regular [Gla89, 6.2.3]. ∎
@@ 4. K-theory and G-theory of semi-valuation rings
${\operatorname{K}(A^{+}[t_{1},\ldots,t_{k}])\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{K}(A[t_{1},\ldots,t_{k}])\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{K}(V[t_{1},\ldots,t_{k}])\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{K}(k[t_{1},\ldots,t_{k}])}$
of non-connective algebraic K-theory spectra is cartesian for every $k\in\mathbf{N}$.
@@
Let $\mathbf{F}$ be the prime field of $k$ so that the valuation ring $V\cap\mathbf{F}$ has rank $\leq 1$. Let $I$ be the set of all subextensions $\ell\subset k$ that have finite transcendence degree over $\mathbf{F}$. Note that $k\cong\operatorname*{colim}_{\ell\in I}\ell$ within the category of rings. Now let $\ell\in I$. By the “dimension inequality” [EP05, Thm. 3.4.3, Cor. 3.4.4], the valuation ring $V_{\ell}:=V\cap\ell$ has finite rank. Hence $V$ is a filtered colimit of valuation rings of finite rank. Set $A_{\ell}$ to be the preimage of $\ell$ in $A$ which is a local ring with residue field $\ell$. Hence the preimage $A^{+}_{\ell}$ of $V$ in $A_{\ell}$ is a semi-valuation ring and the square ()($\square$)( □ ) is the filtered colimit of its restrictions along $\ell\overset{}{\hookrightarrow}k$ for all $\ell\in I$.
Since K-theory commues with filtered colimits of rings we may assume that $V$ has finite rank, hence it is a microbial valuation ring (i.e. has a prime ideal of height 1). In this case, there exists an element $s\in A^{+}$ such that $V[\bar{s}^{-1}]=K$ where $\bar{s}=\pi(s)$. Then one checks easily that $s\in A^{\times}$, $A^{+}[s^{-1}]=A$, and $\mathfrak{p}\subseteq s^{n}\cdot A^{+}$ for all $n\in\mathbf{N}$. Hence the map $\pi\colon A^{+}\overset{}{\rightarrow}V$ is an analytic isomorphism along $S=\{s^{n}\,|\,n\in\mathbf{N}\}$ and the same holds true for the induced maps $A^{+}[t_{1},\ldots,t_{k}]\overset{}{\rightarrow}V[t_{1},\ldots,t_{k}]$ for all $k\in\mathbf{N}$. By Weibel’s analytic isomorphism theorem [Wei80, Thm. 1.3] the square of Theorem 4.1 is cartesian. ∎
@@
where $\operatorname{K}$ denotes Schlichting’s non-connective K-theory for exact categories [Sch04]; cf. Thomason-Trobaugh [TT90, 3.11.1] and Weibel’s K-book [Wei13, V.2.7.4]; note that this can also be realised as the K-theory of a stable $\infty$-category, see [HR19, §8]. If a ring $A$ is regular, then the canonical map $\operatorname{K}(A)\overset{}{\rightarrow}\operatorname{G}(A)$ is an equivalence, since – by definition of regularity – the inclusion $\mathrm{Vec}(A)\overset{}{\hookrightarrow}\mathrm{PCoh}(A)$ satisfies the conditions of the resolution theorem [Wei13, V.3.1].
@@ 5. K-theory of relative Riemann-Zariski spaces
is cartesian in $\mathbf{Sp}$. For $V_{1}^{\prime}\in\mathrm{Model}(V_{1})$ and $V_{2}^{\prime}\in\mathrm{Model}(V_{2})$ we may assume that they live on a common $Y$-modification of $X$. Then $V_{1}^{\prime}\cup V_{2}^{\prime}\in\mathrm{Model}(V)$ and $V_{1}^{\prime}\cap V_{2}^{\prime}\in\mathrm{Model}(V_{3})$. By cofinality, the square ($\spadesuit$) is equivalent to a colimit of cartesian squares since K-theory is a Zariski-sheaf on the category of schemes. The statement for $\operatorname{KH}^{\mathrm{ZR}}(-)$ has the same proof. Since the quasi-compact open subsets of $\operatorname{RZ}_{Y}(X)$ form a basis of the topology, the sheaves extend from $\mathrm{Open}^{\mathrm{qc}}(\operatorname{RZ}_{Y}(X))$ to $\mathrm{Open}(\operatorname{RZ}_{Y}(X))$. ∎
@@
Assume that $Y$ is of finite dimension, that all its stalks are stably coherent regular rings, and that the morphism $f\colon Y\overset{}{\rightarrow}X$ admits a compactification (e.g. $f$ is separated of finite type). Then the properties (i) and (ii) of Proposition 5.9 hold true.

### v3

@@ 1. Introduction
Semi-valuation rings have been introduced by Temkin [Tem11] and they occur as the stalks of his relative Riemann-Zariski spaces $\operatorname{RZ}_{Y}(X)$ which are locally ringed spaces associated with any separated morphism $f\colon Y\overset{}{\rightarrow}X$ between quasi-compact and quasi-separated (qcqs) schemes. Furthermore, semi-valuation rings are important since they are the stalks of discretely ringed adic spaces, see Remark 2.5. Exploiting that algebraic K-theory commutes with the formation of stalks, we deduce the following statement for the K-theory sheaves $\operatorname{K}^{\mathrm{RZ}}$ and $\operatorname{KH}^{\mathrm{RZ}}$ on this space, see Definition 5.6 for a precise definition.
@@
Let $f\colon Y\overset{}{\rightarrow}X$ be a separated morphism between qcqs schemes. Assume that $Y$ is of finite dimension, that all its stalks are stably coherent regular rings, and that the morphism $f\colon Y\overset{}{\rightarrow}X$ admits a compactification (e.g. $f$ is of finite type). Then the canonical maps
@@
Notation. Discrete categories are denoted by $\mathrm{upright}$ $\mathrm{letters}$ whereas genuine $\infty$-categories are denoted by $\mathbf{bold}$ $\mathbf{letters}$. For a ring $R$ we denote by $\operatorname{K}(R)$ its non-connective algebraic K-theory spectrum à la Blumberg-Gepner-Tabuada [BGT13, §9.1] which is an object of the $\infty$-category $\mathbf{Sp}$ of spectra [Lur17, 1.4.3.1]; its associated object in the homotopy category of spectra is equivalent to the K-theory spectrum of Thomason-Trobaugh [TT90, Def. 3.1]. In particular, the connective part $\operatorname{K}_{\geq 0}(R)$ is equivalent to Quillen’s K-theory [Qui73], combine [TT90, Prop. 3.10] with [BGT13, §7.2].
@@ 2. Semi-valuation rings
For the convenience of the reader we recollect the definition and some basic facts about semi-valuation rings. All this is due to Temkin [Tem11]. Let us repeat some relevant terminology: A valuation on a ring $R$ is a map $|-|\colon R\overset{}{\rightarrow}\Gamma\cup\{0\}$ for a totally ordered multiplicative abelian group $\Gamma$ such that $|1|=1$, $|xy|=|x|\cdot|y|$, and $|x+y|\leq\max\{|x|,|y|\}$ for all $x,y\in R$ with the convention that $0\cdot\gamma=0$ and $0<\gamma$ for all $\gamma\in\Gamma$. It follows that $|0|=0$ and that the support $\mathrm{supp}(|-|):=|-|^{-1}(0)$ is a prime ideal. Two valuations $|-|_{1}$ and $|-|_{2}$ on $R$ are called equivalent if for all $x,y\in R$ the conditions $|x|_{1}\leq|y|_{1}$ and $|x|_{2}\leq|y|_{2}$ are equivalent.
@@
A semi-valuation ring is a pair $(A^{+},|-|)$ consisting of a ring $A^{+}$ and a valuation $|-|\colon A^{+}\overset{}{\rightarrow}\Gamma\cup\{0\}$ such that
@@
    every zero divisor of $A^{+}$ lies in the kernel of |||-| and
@@
If $A^{+}$ is a semi-valuation ring and $\mathfrak{p}:=\mathrm{supp}(|-|)$, then the local ring $A^{+}_{\mathfrak{p}}$ is called its semi-fraction ring and the ring $A^{+}/\mathfrak{p}$ is a valuation ring. We call $\mathfrak{p}$ the valuative ideal of $A^{+}$.
@@
    A semi-valuation ring $(A^{+},|-|)$ up to equivalence of the valuation |||-|.
@@
The implications (i) $\Rightarrow$ (ii) $\Leftrightarrow$ (iii) are clear. Assuming (ii), we obtain a valuation $|-|\colon A^{+}\overset{\pi}{\rightarrow}V\overset{}{\rightarrow}V^{\times}/k^{\times}\cup\{0\}$ since $V$ is assumed to be a valuation ring. We will check the conditions (i) and (ii) of Definition 2.1. As $V$ is a domain, we get (i). Now let $x,y\in A^{+}$ with $|x|\leq|y|\neq 0$, hence $y\notin\mathfrak{p}$. Since the morphism $A\overset{}{\rightarrow}k$ reflects units, we get that $y\in A^{\times}$ and that for every $p\in\mathfrak{p}$ the ratio $\frac{p}{y}\in A$ lies in $A^{+}$. By standard properties of valuation rings we find an element $z\in A^{+}$ such that $\pi(x)=\pi(y)\cdot\pi(z)$ in $V$, hence $x=yz+p$ for a suitable element $p\in\mathfrak{p}$. It follows that $x=yz+p=y(z+\frac{p}{y})$ in $A^{+}$ as desired. ∎
@@
In the sequel, we refer to a semi-valuation ring as a pair $(A^{+},\mathfrak{p})$ as in Lemma 2.2, even though the valuation is only defined up to equivalence. We always assume the square $(\square)$ to be implicitly defined.
@@ 3. Coherence and regularity of semi-valuation rings
    $n$-regular if every finitely $n$-presented AA-module has finite projective dimension.
@@
[…]
@@
The notion “1-regular” is called “regular” by Gersten [Ger74, Def. 1.3]. The notion “Glaz-regular” is called “regular” by Glaz [Gla89, ch. 6, §2]. For a coherent ring, these notions agree with Waldhausen’s notion “regular coherent” [Wal78, p. 138]. Antieau-Mathew-Morrow [AMM22, §2] call a ring “weakly regular” if it has finite flat dimension (for them a regular ring is noetherian). For coherent rings, the notion “weakly regular” is equivalent to the notion “uniformly 1-regular” and it is stronger than the notions used by Gersten and Glaz. For arbitrary rings, the notion “regular” from Definition 3.2 seems to be most meaningful (at least to the author).
@@
(i) Let $x_{1},\ldots,x_{n}\in M$ such that their images form a $R/\mathfrak{m}$-basis of $M/\mathfrak{m}M$ so that the morphism $\varphi\colon R^{n}\overset{}{\rightarrow}M,e_{i}\mapsto x_{i},$ induces an isomorphism $\bar{\varphi}\colon(R/\mathfrak{m})^{n}\overset{}{\rightarrow}M/\mathfrak{m}M$. By Nakayama’s lemma we get an exact sequence $0\overset{}{\rightarrow}N\overset{}{\rightarrow}R^{n}\overset{\varphi}{\rightarrow}M\overset{}{\rightarrow}0$, inducing an exact sequence
$0=\operatorname{Tor}_{1}^{R}(M,R/\mathfrak{m})\overset{}{\longrightarrow}N/\mathfrak{m}N\overset{}{\longrightarrow}(R/\mathfrak{m})^{n}\overset{\cong}{\longrightarrow}M/\mathfrak{m}M\overset{}{\longrightarrow}0$
@@
be an exact sequence where $F_{n+1},\ldots,F_{0}$ are finitely generated free. Setting $K_{i}:=\ker(f_{i})$ we get that
@@
[…]
@@
By Lemma 2.4, we can write $\mathfrak{m}=\operatorname*{colim}_{x\in I}(x)$ as a filtered colimit with $I=\mathfrak{m}^{+}\setminus\mathfrak{p}\neq\emptyset$. According to condition (i) in Definition 2.1, the sequence $0\overset{}{\rightarrow}A^{+}\overset{\cdot x}{\rightarrow}A^{+}\overset{}{\rightarrow}A^{+}/(x)\overset{}{\rightarrow}0$ is exact for every $x\in I$. Since tensor products are cocontinuous in each variable and since filtered colimits are exact, we get that
@@
Given a Milnor square (M) such that the morphism $A\overset{f}{\rightarrow}B$ is of finite Tor-dimension and assuming that the rings $A^{\prime}$, $B$, and $B^{\prime}$ all are regular. Is then the ring $A$ regular?
@@
[…]
@@
We know that the ring $A^{+}$ is 2-regular by Lemma 3.7. If it was also coherent, then it would be Glaz-regular by Proposition 3.3, hence AA would be Glaz-regular [Gla89, 6.2.3]. ∎
@@ 4. K-theory and G-theory of semi-valuation rings
${\operatorname{K}(A^{+}[t_{1},\ldots,t_{r}])\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{K}(A[t_{1},\ldots,t_{r}])\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{K}(V[t_{1},\ldots,t_{r}])\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{K}(k[t_{1},\ldots,t_{r}])}$
of non-connective algebraic K-theory spectra is cartesian for every $r\in\mathbf{N}$.
@@
Let $\mathbf{F}$ be the prime field of $k$ so that the valuation ring $V\cap\mathbf{F}$ has rank $\leq 1$. Let $I$ be the set of all subextensions $\ell\subset k$ that have finite transcendence degree over $\mathbf{F}$. Note that $k\cong\operatorname*{colim}_{\ell\in I}\ell$ within the category of rings. Now let $\ell\in I$. By the “dimension inequality” [EP05, Thm. 3.4.3, Cor. 3.4.4], the valuation ring $V_{\ell}:=V\cap\ell$ has finite rank. Hence $V$ is a filtered colimit of valuation rings of finite rank. Set $A_{\ell}$ to be the preimage of $\ell$ in $A$ which is a local ring with residue field $\ell$. Hence the preimage $A^{+}_{\ell}$ of $V$ in $A_{\ell}$ is a semi-valuation ring and the square $(\square)$ is the filtered colimit of its restrictions along $\ell\overset{}{\hookrightarrow}k$ for all $\ell\in I$.
Since K-theory commutes with filtered colimits of rings we may assume that $V$ has finite rank, hence it is a microbial valuation ring (i.e. has a prime ideal of height 1). In this case, there exists an element $s\in A^{+}$ such that $V[\bar{s}^{-1}]=K$ where $\bar{s}=\pi(s)$ for the projection $\pi\colon A^{+}\overset{}{\rightarrow}V$. Then one checks easily that $s\in A^{\times}$, $A^{+}[s^{-1}]=A$, and $\mathfrak{p}\subseteq s^{n}\cdot A^{+}$ for all $n\in\mathbf{N}$. Hence the map $\pi\colon A^{+}\overset{}{\rightarrow}V$ is an analytic isomorphism along $S=\{s^{n}\,|\,n\in\mathbf{N}\}$ and the same holds true for the induced maps $A^{+}[t_{1},\ldots,t_{r}]\overset{}{\rightarrow}V[t_{1},\ldots,t_{r}]$ for all $r\in\mathbf{N}$. By Weibel’s analytic isomorphism theorem [Wei80, Thm. 1.3] the square of Theorem 4.1 is cartesian. Note that every Milnor square induces a cartesian square on non-positive K-theory $\operatorname{K}_{\leq 0}$ [Bas68, Ch. XII, Thm. (8.3), p. 677]. ∎
@@
where $\operatorname{K}$ denotes Schlichting’s non-connective K-theory for exact categories [Sch04]; c.f. Thomason-Trobaugh [TT90, 3.11.1] and Weibel’s K-book [Wei13, V.2.7.4]; note that this can also be realised as the K-theory of a stable $\infty$-category, see [HR19, §8]. If a ring $A$ is regular, then the canonical map $\operatorname{K}(A)\overset{}{\rightarrow}\operatorname{G}(A)$ is an equivalence, since – by definition of regularity – the inclusion $\mathrm{Vec}(A)\overset{}{\hookrightarrow}\mathrm{PCoh}(A)$ satisfies the conditions of the resolution theorem [Wei13, V.3.1].
@@ 5. K-theory of relative Riemann-Zariski spaces
is cartesian in $\mathbf{Sp}$. For $V_{1}^{\prime}\in\mathrm{Model}(V_{1})$ and $V_{2}^{\prime}\in\mathrm{Model}(V_{2})$ we may assume that they live on a common $Y$-modification of $X$. Then $V_{1}^{\prime}\cup V_{2}^{\prime}\in\mathrm{Model}(V)$ and $V_{1}^{\prime}\cap V_{2}^{\prime}\in\mathrm{Model}(V_{3})$. By cofinality, the square ($\spadesuit$) is equivalent to a colimit of cartesian squares since K-theory is a Zariski-sheaf on the category of qcqs schemes. The statement for $\operatorname{KH}^{\mathrm{ZR}}(-)$ has the same proof. Since the quasi-compact open subsets of $\operatorname{RZ}_{Y}(X)$ form a basis of the topology, the sheaves extend from $\mathrm{Open}^{\mathrm{qc}}(\operatorname{RZ}_{Y}(X))$ to $\mathrm{Open}(\operatorname{RZ}_{Y}(X))$. ∎
@@
Assume that $Y$ is of finite dimension, that all its stalks are stably coherent regular rings, and that the morphism $f\colon Y\overset{}{\rightarrow}X$ admits a compactification (e.g. $f$ is of finite type). Then the properties (i) and (ii) of Proposition 5.9 hold true.